Find the value of x.
A triangle. One exterior angle is 70 and the nonadjacent interior angles are X and 2X + 10.

Answer by stanbon(75887)

A triangle. One exterior angle is 70 and the nonadjacent interior angles are X and 2X + 10.
---
The sum of the exterior angles is 360 degrees.
The exterior corresponding to "x" is 180-x
The exterior correspondiing to "2x+10" = 180-(2x+10) = -2x+170
---
Solve:
(180-x)+(-2x+170) + 70 = 360
-3x + 420 = 360
-3x = -60
x = 20 degrees
=================
Cheers,
Stan H.
============
===========================